Please can someone help me with this problem:

Algebra ->  Coordinate Systems and Linear Equations -> SOLUTION: "Tutor" opeodu needs a real tutor because he cannot solve the problem (see below). Hint: answer will be (0, -3): Question 624942: Please can someone help me with this problem:       Log On


   



Question 625125: "Tutor" opeodu needs a real tutor because he cannot solve the problem (see below). Hint: answer will be (0, -3):
Question 624942: Please can someone help me with this problem:
Solve using the elimination method.
14x + 18y = -54
-4x - 14y = 42
Answer by opeodu(4)

14x+18y=-54------(a)*(-4)
-4x-14y=42-------(b)*(14)
-56x-72y=216
-56x-196y=588
subtract
-72y-(-196y)=216-588
124y=372
y=372/124=3
substitute y=3 in (a)
14x+18y=-54
14x+18(3)=-54
14x=-54-54
14x=-108
x=-108/14=7whole number5 over 7


Answer by oscargut(2103)

Hi,
Yes, the answer above is wrong
The answer is:
x = 0 and y = -3
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com